OpenStudy (anonymous):

Given that the Moon subtends an angle of about 0.5 deg, and has a diameter of 3476 km, find the distance to the Moon using the small angle formula.

OpenStudy (anonymous):

\[c = \pi d\]

OpenStudy (anonymous):

c = 22/7 \[\times \] d = 3476

OpenStudy (anonymous):

so now u can proceed...